# An Iconic Heavy Hitter Algorithm Made Private

### v1

@@ 3.2 Linear Sketches
Linear sketches enable approximate frequency queries, in small memory, by hashing items to a compact collection of shared counters. We focus on the well known Count-min Sketch ($\mathsf{CountMinSketch}$) cormode2005improved and Count Sketch ($\mathsf{CountSketch}$) charikar2002finding variants. While they share a similar structure, they differ in their update and query procedures and offer different error trade-offs.
@@ 3.2.1 Count-Min Sketch
The $\mathsf{CountMinSketch}$ uses $d$ hash functions $h_{1},\ldots,h_{d}:\mathcal{U}\rightarrow[w]$ to maintain a $d\times w$ matrix of counters $C\in(\mathrm{R}^{+})^{d\times w}$. For each incoming item $x\in\mathcal{U}$ the update rule is:
@@
Lemma 4 (Approximation error for $\mathsf{CountMinSketch}$ cormode2005improved).
For $x\in\mathcal{U}$ and any $\beta,\eta\in(0,1)$, the estimation error of Count Min Sketch with $d=\lceil\log(1/\beta)\rceil$ repetitions and table size $w=\lceil 2/\eta\rceil$ satisfies,
$\mathds{P}[\hat{f}_{x}-f_{x}>\eta F_{1}]\leq\beta,$
where $F_{1}=\sum_{y\in\mathcal{U}}f_{y}$.
@@ 3.2.2 Count Sketch
Lemma 5 (Approximation error for $\mathsf{CountSketch}$ charikar2002finding).
$\Pr\left[|\hat{f}_{x}-f_{x}|>\eta\sqrt{F_{2}/w}\right]\leq\beta,$
where $F_{2}=\sum_{y\in[n]}f_{y}^{2}$ is the second moment of the frequency vector.
4 Differentially Private SpaceSaving
@@
To privatize a counter-based algorithm it is not sufficient to inject noise into the counts of the tracked items. This is because the IDs of the tracked items can differ on neighboring inputs, allowing an adversary to easily distinguish the streams based on the output of the algorithm. To overcome this challenge for $\mathsf{MisraGries}$, Lebeda and Tetek enumerate all possible differences between the structures $\mathsf{MG}(X,k)$ and $\mathsf{MG}(X^{\prime},k)$, for $X\sim X^{\prime}$. Knowledge of these differences allows a non-private $\mathsf{MisraGries}$ to be processed such that both privacy leakage events are suppressed and item count differences are obscured. We adopt the same approach for privatizing $\mathsf{SpaceSaving}$.
@@ 4.1 SpaceSaving on Neighboring Streams
Before proceeding, we provide some useful notation. Let $X=\{x_{1},\ldots,x_{T}\}$ and $X^{\prime}=\{x_{1},\ldots,x_{i-1},\varnothing,x_{i+1},\ldots,x_{T}\}$ denote neighboring inputs, let $W_{t}=\mathcal{T}_{t}\setminus\mathcal{T}_{t}^{\prime}$ and $W_{t}^{\prime}=\mathcal{T}_{t}^{\prime}\setminus\mathcal{T}_{t}$ denote the set difference functions. Item eviction is determined by the recency functions $\tau_{t}[y]=\max\{r\mid X[r]=y,\;r\leq t\}$ and $\tau_{t}^{\prime}[y]=\max\{r\mid X^{\prime}[r]=y,\;r\leq t\}$. Define the sets of items with minimum count as $S_{i+1}=\arg\min\{C_{i}[z]\mid z\in\mathcal{T}_{i}\}$ and $S_{i+1}^{\prime}=\arg\min\{C_{i}^{\prime}[z]\mid z\in\mathcal{T}_{i}^{\prime}\}$. Thus, our eviction strategy at time $t$ is $\arg\max_{a\in S_{t}}\tau_{t-1}[a]$. We refer to this item as occupying the eviction register.
To formalize the differences between the outputs of the $\mathsf{SpaceSaving}$ algorithm on neighboring streams, we categorize each type of difference into a state. Our main privacy proof requires the enumeration of all possible states, which are provided in Table 1 for reference. This phenomenon is summarized by the following result.
@@
Let $\mathsf{SS}(X,k)=(\mathcal{T}_{t},C_{t})$ and $\mathsf{SS}(X^{\prime},k)=(\mathcal{T}_{t}^{\prime},C_{t}^{\prime})$ be the outputs of Algorithm 2 on neighboring inputs $X$ and $X^{\prime}$ such that $X^{\prime}$ is obtained by removing an element from $X$. Then, $|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq k-2$ and all counters not in the intersection have value at most $\min_{a\in\mathcal{T}^{\prime}}C^{\prime}[a]+1$. In addition, there exists at most one item $x\in\mathcal{T}\cap\mathcal{T}^{\prime}$ such that $C[x]=C^{\prime}[x]+1$ and $\forall y\in\mathcal{T}\cap\mathcal{T}^{\prime}\setminus\{x\}:C[y]=C[y^{\prime}]$.
The proof of Lemma 6 proceeds inductively. We break the proof up into a sequence of Lemmas for readability and begin with the base case (time $t=i$).
@@
Figure 1: Example of transition to $(\mathsf{S1})$ when element $z\notin\mathcal{T}$ arrives from $X$ and $X^{\prime}$ provides no update.
@@
There are two cases to consider for update $i$. First, we have $x_{i}\in\mathcal{T}_{i-1}$. In this instance the counter $C_{i-1}[x_{i}]$ gets incremented to form $C_{i}[x_{i}]$ and the set of tracked items stays the same (condition $(\mathsf{S1})$(a)). As $\mathcal{T}_{i-1}^{\prime}=\mathcal{T}_{i}$ and $C_{i-1}^{\prime}=C_{i}^{\prime}$, conditions $(\mathsf{S1})$(b) and $(\mathsf{S1})$(c) are satisfied.
@@
Examples of the possible state transitions at time $i$ are given in Figures 1 and 2. Before proceeding with the state differences at times $t>i$, we prove a simple but useful result.
@@
The proof is relegated to Appendix B, due to its similarity to the proof of Lemma 9. Examples of state transitions from $(\mathsf{S2})$ are given in Figures 45. The transition to $(\mathsf{S1})$ occurs when the isolated elements of both structures are in the eviction register. In contrast, the transition to $(\mathsf{S3})$ occurs when the isolated element in $\mathcal{T}_{t-1}$ is not in the eviction register. Both transitions can be triggered when an item $y\notin\mathcal{T}_{t-1}\cup\mathcal{T}_{t-1}^{\prime}$ arrives.
@@
The proof is relegated to Appendix B. An example of a transition from $(\mathsf{S3})$ to $(\mathsf{S1})$ is given in Figure 6. The transition is triggered if an isolated item is in the eviction register of $\mathcal{T}_{t-1}$. An example of a transition from $(\mathsf{S3})$ to $(\mathsf{S4})$ is given in Figure 7 and increases the number of isolated elements. The transition is triggered if the eviction registers are not identical and do not contain isolated elements. This occurs when the eviction register in $\mathcal{T}_{t-1}^{\prime}$ is occupied by the unique element $y\in\mathcal{T}_{t-1}\cap\mathcal{T}_{t-1}^{\prime}$ where $C_{t-1}[y]=C_{t-1}^{\prime}[y]+1$.
@@
The proof is relegated to Appendix B. An important condition of $(\mathsf{S4})$ is that isolated elements from both streams are always in the eviction registers. This restricts available transitions to $(\mathsf{S2})$ or $(\mathsf{S3})$ only, as, if an element outside the intersection $\mathcal{T}_{t-1}\cap\mathcal{T}_{t-1}^{\prime}$ arrives, the number of isolated elements must decrease. The resulting state depends on the identify of the element and examples are given in Figures 89.
@@
Proceeding inductively, Lemma 7 provides the base case and Lemmas 9-12 provide the subsequent inductive steps. ∎
@@ 4.2 Privacy
Second, it guarantees that items outside the intersection are associated with low counts, specifically bounded above by the minimum counter in $\mathcal{T}^{\prime}$ plus one. These properties together imply that the output of $\mathsf{SpaceSaving}$ is inherently stable, and thus amenable to privatization via two mechanisms: (1) obscuring the small count differences among shared items, and (2) suppressing isolated items whose counts fall below a threshold.
The first mechanism is realized by injecting Laplace noise into the counters of tracked items. Since the sensitivity of the counters restricted to $\mathcal{T}\cap\mathcal{T}^{\prime}$ is at most one, Lemma 1 ensures that adding independent noise drawn from $\mathtt{Laplace}(1/\varepsilon)$ suffices to achieve differential privacy.
The second step can be achieved by applying a threshold test that suppresses items with low count. Let $\gamma$ denote the maximum plausible deviation introduced by $\mathtt{Laplace}$ noise. In $\mathsf{SpaceSaving}$, for a register of length $k$, the minimum count is at most $|X|/k$. Therefore, it would be sufficient to set the threshold to $|X|/k+1+\gamma$. However, as our task is to detect items whose frequency is at least $|X|/k$ (heavy hitters), this introduces a tension. Notably, overly aggressive suppression risks excluding valid heavy hitters that are permissible to report.
To reconcile privacy with utility, we initialize a non-private $\mathsf{SpaceSaving}$ data structure with an expanded capacity $\tilde{k}>k$. This adjustment ensures that items with minimal counts are unlikely to be heavy hitters, thereby allowing suppression without sacrificing recall. This strategy is formalized in Algorithm 3. We now prove the privacy properties.
@@ 4 $\hat{\mathcal{T}},\hat{C}\leftarrow\varnothing$
5 $\tau\leftarrow\max\{|X|/k,|X|/\tilde{k}+1+\gamma\}$
@@ 10       $\hat{C}[x]\leftarrow\hat{c}_{x}$
where $Z_{u}$ is the Laplace noise. As $\mathsf{SpaceSaving}$ is composed of $\tilde{k}$ counters that sum to $|X|$, the minimum count is at most $|X|/\tilde{k}$. Therefore, by Corollary 1 and Lemma 3, $|U|\leq 2$ and each such $u\in U$ has true count at most $|X|/\tilde{k}+1$. Since $\tau\geq|X|/\tilde{k}+1+\gamma$, the margin from the minimum count to the threshold is at least $\gamma$ for every such $u\in U$. For Laplace noise $Z\sim\mathrm{Lap}(1/\varepsilon)$ we have the standard tail bound $\Pr[\,Z\geq\gamma\,]=\tfrac{1}{2}e^{-\varepsilon\gamma}.$ Applying a union bound over at most two unstable labels and two independent runs yields
@@ 4.3 Performance and Utility
Algorithm 3 inherits the efficiency of the underlying $\mathsf{SpaceSaving}$ algorithm, offering strong performance guarantees in both time and space. Using an implementation based on hash tables and doubly linked lists Algorithm 3 can be implemented with O(k~)O($\tilde${k}) space complexity and can process each stream update in $O(1)$ time cormode2008finding. This ensures that the algorithm scales gracefully with high-throughput data streams.
In practice, for large streams, the minimum counter value $|X|/\tilde{k}$ dominates the noise parameter $\gamma=\frac{1}{\varepsilon}\ln\left(\frac{2}{\delta}\right)$. Consequently, it is sufficient to set $\tilde{k}=O(k)$ without compromising utility. Thus, the additional space does not come at an asymptotical cost.
@@
$\tau\;=\;\max\Bigl\{\,|X|/k,\;|X|/\tilde{k}+1+\gamma\,\Bigr\},\quad\text{where}\quad\gamma\;=\;\frac{1}{\varepsilon}\ln\!\left(\frac{2}{\delta}\right).$
To ensure $\tau=|X|/k$, so heavy hitters are not suppressed, we require
$\frac{|X|}{\tilde{k}}+1+\gamma\leq\frac{|X|}{k}\quad\Rightarrow\quad\tilde{k}\geq\frac{|X|}{\frac{|X|}{k}-(1+\gamma)}.$
@@
For any $\varepsilon>0$ and $\delta\in(0,1)$, let $\gamma=\varepsilon^{-1}\ln(2\delta^{-1})$. Suppose the input stream $X$ and target capacity $k$ satisfy $|X|/(2k)>\gamma\;$. Then Algorithm 3 can be implemented with $\mathcal{O}(k)$ words of memory and $\mathcal{O}(1)$ update time. Moreover, for any item $x\in\hat{\mathcal{T}}$ with true frequency $f_{x}$, with probability at least $1-\delta$ the released estimate $\hat{C}[x]$ obeys
$f_{x}\;-\;\mathcal{O}\!\left(\tfrac{1}{\varepsilon}\ln\tfrac{1}{\delta}\right)\;\leq\;\hat{C}[x]\;\leq\;f_{x}+\frac{|X|}{2k}+\mathcal{O}\!\left(\tfrac{1}{\varepsilon}\ln\tfrac{1}{\delta}\right).$
@@
Proof.
Set $\tilde{k}=2k$ and run $\mathsf{SpaceSaving}$ with capacity $\tilde{k}$. This uses $\mathcal{O}(\tilde{k})=\mathcal{O}(k)$ words and $\mathcal{O}(1)$ update time. Let $C[\cdot]$ be the terminal (non-private) counters and add independent noise $\eta_{x}\sim\mathrm{Laplace}(1/\varepsilon)$. Release only labels with $C[x]+\eta_{x}>\tau$, where
$\tau\;=\;\frac{|X|}{\tilde{k}}\;+\;\gamma\;=\;\frac{|X|}{2k}\;+\;\frac{1}{\varepsilon}\ln\!\Big(\frac{2}{\delta}\Big).$
By deterministic guarantee of $\mathsf{SpaceSaving}$ (Lemma 3),
$f_{x}\;\leq\;C[x]\;\leq\;f_{x}+\frac{|X|}{2k}\qquad\text{for all items }x.$
For $\eta_{x}\sim\mathrm{Laplace}(1/\varepsilon)$ we have $\Pr[|\eta_{x}|>t]=e^{-\varepsilon t}$. Taking $t=\frac{1}{\varepsilon}\ln(1/\delta)$ gives, with probability $\geq 1-\delta$,
which matches the stated bounds.
If $f_{y}>|X|/k$, then $C[y]\geq f_{y}\geq|X|/k$, so
$C[y]-\tau\;\geq\;\frac{|X|}{k}-\Big(\frac{|X|}{2k}+\gamma\Big)=\frac{|X|}{2k}-\gamma\;>\;0$
by the premise $|X|/(2k)>\gamma$. Therefore $y$ fails the threshold only if $\eta_{y}<-(|X|/(2k)-\gamma)$, which occurs with probability
$\leq\tfrac{1}{2}e^{-\varepsilon(|X|/(2k)-\gamma)}\leq\delta/2.$
Combining with the suppression bound yields $y\in\hat{\mathcal{T}}$ with probability $\geq 1-\delta$. ∎

### v2

@@ 3.2 Linear Sketches
Linear sketches enable approximate frequency queries, in small memory, by hashing items to a compact collection of shared counters. We focus on the well known Count-min Sketch ($\mathsf{CountMinSketch}$Cormode and Muthukrishnan (2005) and Count Sketch ($\mathsf{CountSketch}$Charikar et al. (2002) variants. While they share a similar structure, they differ in their update and query procedures and offer different error trade-offs.
@@ 3.2.1 Count-Min Sketch
The $\mathsf{CountMinSketch}$ uses $d$ hash functions $h_{1},\ldots,h_{d}:\mathcal{U}\rightarrow[w]$ to maintain a $d\times w$ matrix of counters $C\in(\mathbb{R}^{+})^{d\times w}$. For each incoming item $x\in\mathcal{U}$ the update rule is:
@@ 3.2.2 Count Sketch
3.2.3 $\mathsf{DP}$ Heavy Hitters with $\mathsf{DP}$ Linear Sketches
@@
A $\mathsf{DP}$ linear sketch can be used as a frequency oracle within a heavy hitter search algorithm. The most direct approach performs a brute-force scan over the entire item universe Epasto et al. (2023), which is computationally intractable for large (which is to say data stream) domains. More efficient methods employ tree-based search procedures that recursively prune infrequent regions of the domain Bassily et al. (2017). Nevertheless, these approaches still require $\mathcal{O}(\sqrt{d})$ oracle queries, where $d$ denotes an upper bound on the number of distinct items, resulting in non-trivial (and non-streaming) computational and memory overheads.
Alternatively, in the non-private setting, heavy hitters can be extracted from a linear sketch by tracking the items with the largest observed frequency estimates throughout the stream. However, as with counter-based algorithms, the resulting candidate sets may differ across neighboring streams, rendering their direct release incompatible with $\mathsf{DP}$. Designing a privacy-preserving analogue of this tracking-based approach is non-trivial, and addressing this challenge is a central contribution of our work.
4 Differentially Private SpaceSaving
To privatize a counter-based algorithm it is not sufficient to inject noise into the counts of the tracked items. This is because the identities of the tracked items can differ on neighboring inputs, allowing an adversary to easily distinguish the streams based on the output of the algorithm. To overcome this challenge for $\mathsf{MisraGries}$, Lebeda and Tetek enumerate all possible differences between the structures $\mathsf{MG}(X,k)$ and $\mathsf{MG}(X^{\prime},k)$, for $X\sim X^{\prime}$ Lebeda and Tetek (2023). Knowledge of these differences allows a non-private $\mathsf{MisraGries}$ to be processed such that both privacy leakage events are suppressed and item count differences are obscured. We adopt the same approach for privatizing $\mathsf{SpaceSaving}$.
@@ 4.1 SpaceSaving on Neighboring Streams
Before proceeding, we provide some useful notation. Let $X=\{x_{1},\ldots,x_{T}\}$ and $X^{\prime}=\{x_{1},\ldots,x_{i-1},\varnothing,x_{i+1},\ldots,x_{T}\}$ denote neighboring inputs, let $(\mathcal{T}_{t},C_{t})\leftarrow\mathsf{SS}(X_{t})$ denote the output of $\mathsf{SpaceSaving}$ at time $t$, and let $W_{t}=\mathcal{T}_{t}\setminus\mathcal{T}_{t}^{\prime}$ and $W_{t}^{\prime}=\mathcal{T}_{t}^{\prime}\setminus\mathcal{T}_{t}$ denote the set difference functions. Item eviction is determined by the recency functions $\tau_{t}[y]=\max\{r\mid X[r]=y,\;r\leq t\}$ and $\tau_{t}^{\prime}[y]=\max\{r\mid X^{\prime}[r]=y,\;r\leq t\}$. Define the sets of items with minimum count, prior to eviction, as $S_{t+1}=\arg\min\{C_{t}[z]\mid z\in\mathcal{T}_{t}\}$ and $S_{t+1}^{\prime}=\arg\min\{C_{t}^{\prime}[z]\mid z\in\mathcal{T}_{t}^{\prime}\}$. Thus, our eviction strategy at time $t$ is $\arg\max_{a\in S_{t}}\tau_{t-1}[a]$. We refer to this item as occupying the eviction register.
To formalize the differences between the outputs of the $\mathsf{SpaceSaving}$ algorithm on neighboring streams, we categorize each type of difference into a state. Our main privacy proof requires the enumeration of all possible states, which are provided in Table 1. The following lemma establishes that this enumeration is exhaustive.
@@
Let $\mathsf{SS}(X,k)=(\mathcal{T},C)$ and $\mathsf{SS}(X^{\prime},k)=(\mathcal{T}^{\prime},C^{\prime})$ be the outputs of Algorithm 2 on neighboring streams $X$ and $X^{\prime}$, where $X^{\prime}$ is obtained from $X$ by removing a single update. Then the following properties hold:
  1. 1.
    The intersection satisfies $|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq k-2$.
  2. 2.
    For every item $a\in\mathcal{T}\setminus(\mathcal{T}\cap\mathcal{T}^{\prime})$, its counter obeys
    $C[a]\leq\min_{b\in\mathcal{T}}C[b]+1,$
    and for every item $a\in\mathcal{T}^{\prime}\setminus(\mathcal{T}\cap\mathcal{T}^{\prime})$,
    $C^{\prime}[a]\leq\min_{b\in\mathcal{T}^{\prime}}C^{\prime}[b].$
  3. 3.
    There exists at most one item $x\in\mathcal{T}\cap\mathcal{T}^{\prime}$ such that $C[x]=C^{\prime}[x]+1$, and for all $y\in(\mathcal{T}\cap\mathcal{T}^{\prime})\setminus\{x\}$, we have $C[y]=C^{\prime}[y]$.
This corollary informs the design of a $\mathsf{DP}$ post-processing procedure for the output of the non-private $\mathsf{SpaceSaving}$ algorithm. Before outline such a procedure, we turn to the formal justification of this claim by establishing Lemma 4.
The proof of Lemma 4 proceeds by induction on the stream length. We break the proof up into a sequence of intermediate lemmas for readability and begin with the base case (time $t=i$).
@@
Figure 1: Example of transition to $(\mathsf{S1})$ when element $y_{1}\in\mathcal{T}$ arrives from $X$ and $X^{\prime}$ provides no update.
@@
There are two cases to consider for update $i$. First, we have $x_{i}\in\mathcal{T}_{i-1}(=\mathcal{T}_{i-1}^{\prime})$. In this instance the counter $C_{i-1}[x_{i}]$ gets incremented to form $C_{i}[x_{i}]$, and the set of tracked items stays the same (condition $(\mathsf{S1})$(a)). As $C_{i-1}^{\prime}=C_{i}^{\prime}$, conditions $(\mathsf{S1})$(b) and $(\mathsf{S1})$(c) are satisfied.
@@
Examples of the possible state transitions at time $i$ are given in Figures 12. Before proceeding with the state differences at times $t>i$, we prove a simple but useful result.
@@
The proof is relegated to Appendix C, due to its similarity to the proof of Lemma 7. Examples of state transitions from $(\mathsf{S2})$ are given in Figures 45. The transition to $(\mathsf{S1})$ occurs when the isolated elements of both structures are in the eviction register. In contrast, the transition to $(\mathsf{S3})$ occurs when the isolated element in $\mathcal{T}_{t-1}$ is not in the eviction register. Both transitions can be triggered when an item $y\notin\mathcal{T}_{t-1}\cup\mathcal{T}_{t-1}^{\prime}$ arrives.
@@
The proof is relegated to Appendix C. An example of a transition from $(\mathsf{S3})$ to $(\mathsf{S1})$ is given in Figure 6. The transition is triggered if an isolated element is in the eviction register of $\mathcal{T}_{t-1}$ but not $\mathcal{T}_{t-1}^{\prime}$. An example of a transition from $(\mathsf{S3})$ to $(\mathsf{S4})$ is given in Figure 7 and increases the number of isolated elements. The transition is triggered if the eviction registers are not identical and do not contain isolated elements. This occurs when the eviction register in $\mathcal{T}_{t-1}^{\prime}$ is occupied by the unique element $y\in\mathcal{T}_{t-1}\cap\mathcal{T}_{t-1}^{\prime}$ where $C_{t-1}[y]=C_{t-1}^{\prime}[y]+1$.
@@
The proof is relegated to Appendix C. An important condition of $(\mathsf{S4})$ is that isolated elements from both streams are always in the eviction registers. This restricts available transitions to $(\mathsf{S2})$ or $(\mathsf{S3})$ only, as, if an element outside the intersection $\mathcal{T}_{t-1}\cap\mathcal{T}_{t-1}^{\prime}$ arrives, the number of isolated elements must decrease. The resulting state depends on the identify of the element and examples are given in Figures 89.
@@
Proceeding inductively, Lemma 5 provides the base case and Lemmas 7-10 provide the subsequent inductive steps. ∎
@@ 4.2 Privacy
Second, it guarantees that items outside the intersection are associated with low counts, specifically bounded above by the minimum counter plus one. These properties together imply that the output of $\mathsf{SpaceSaving}$ is inherently stable, and thus amenable to privatization via two mechanisms: (1) obscuring the small count differences among shared items, and (2) suppressing isolated items whose counts fall below a threshold.
The first mechanism is realized by injecting $\mathtt{Laplace}$ noise into the counters of tracked items. Since the sensitivity of the counters restricted to $\mathcal{T}\cap\mathcal{T}^{\prime}$ is at most one, Lemma 1 ensures that adding independent noise drawn from $\mathtt{Laplace}(1/\varepsilon)$ suffices to achieve differential privacy.
The second step can be achieved by applying a threshold test that suppresses items with low count. Let $\gamma$ denote, with probability $1-\delta$, the maximum deviation introduced by $\mathtt{Laplace}$ noise. In $\mathsf{SpaceSaving}$, for a register of length $k$, the minimum count is at most $|X|/k$. Therefore, it would be sufficient to set the threshold to $|X|/k+1+\gamma$. However, as our task is to detect items whose frequency is at least $|X|/k$ (heavy hitters), this introduces a tension. Notably, overly aggressive suppression risks excluding valid heavy hitters that can be reported without breaching privacy.
To reconcile privacy with utility, we initialize a non-private $\mathsf{SpaceSaving}$ data structure with an expanded capacity $\tilde{k}>k$. This adjustment ensures that items with minimal counts are unlikely to be heavy hitters, thereby allowing suppression without sacrificing recall. Subsequently, it is sufficient to set the threshold to
$\tau\leftarrow\max\{|X|/k-\gamma,|X|/\tilde{k}+1+\gamma\}.$
With high probability, the term $|X|/k-\gamma$ guarantees the admission of heavy hitters and the term $|X|/\tilde{k}+1+\gamma$ suppresses isolated items. This strategy is formalized in Algorithm 3. We now prove the privacy properties.
@@ 4 $\hat{\mathcal{T}},\hat{C}\leftarrow\varnothing$
5 $\tau\leftarrow\max\{|X|/k-\gamma,|X|/\tilde{k}+1+\gamma\}$
@@ 10       $\hat{C}[x]\leftarrow\hat{c}_{x}$
where $Z_{u}$ is the Laplace noise. As $\mathsf{SpaceSaving}$ is composed of $\tilde{k}$ counters that sum to $|X|$, the minimum count is at most $|X|/\tilde{k}$. Therefore, by Corollary 1 and Lemma 3, $|U|\leq 2$ and each such $u\in U$ has true count at most $|X|/\tilde{k}+1$. Since $\tau\geq|X|/\tilde{k}+1+\gamma$, the margin from the minimum count to the threshold is at least $\gamma$ for every such $u\in U$. For Laplace noise $Z\sim\mathrm{Lap}(1/\varepsilon)$ we have the standard tail bound $\Pr[\,Z>\gamma\,]=\tfrac{1}{2}e^{-\varepsilon\gamma}.$ Applying a union bound over at most two unstable labels and two independent runs yields
@@
Condition on $E$, we have $\mathcal{A}(X)=\Phi(F(X)+Z)$. Putting everything together, using (2), $\Pr[E]\leq 1$, and $\Pr[\neg E]\leq\delta$, we obtain
@@ 4.3 Performance and Utility
Algorithm 3 inherits the efficiency of the underlying $\mathsf{SpaceSaving}$ algorithm, offering strong performance guarantees in both time and space. Using an implementation based on hash tables and doubly linked lists, Algorithm 3 can be implemented with O(k~)O($\tilde${k}) space complexity and can process each stream update in $O(1)$ time Cormode and Hadjieleftheriou (2008). This ensures that the algorithm scales gracefully with high-throughput data streams.
In practice, for large streams, the bound on the minimum counter value $|X|/\tilde{k}$ dominates the noise parameter $\gamma=\frac{1}{\varepsilon}\ln\left(\frac{2}{\delta}\right)$. Consequently, it is sufficient to set $\tilde{k}=O(k)$ without compromising utility. Thus, the additional space does not come at an asymptotical cost.
@@
$\tau\;=\;\max\Bigl\{\,|X|/k-\gamma,\;|X|/\tilde{k}+1+\gamma\,\Bigr\},\quad\text{where}\quad\gamma\;=\;\frac{1}{\varepsilon}\ln\!\left(\frac{2}{\delta}\right).$
To ensure $\tau=|X|/k-\gamma$, so heavy hitters are not suppressed, we require
$\frac{|X|}{\tilde{k}}+1+2\gamma\leq\frac{|X|}{k}\quad\Rightarrow\quad\tilde{k}\geq\frac{|X|}{\frac{|X|}{k}-(1+\gamma)}.$
@@
For any $\varepsilon>0$ and $\delta\in(0,1)$, let $\gamma=\tfrac{1}{\varepsilon}\ln\left(\tfrac{2}{\delta}\right)$. Suppose the input stream $X$ and target capacity $k$ satisfy
$|X|/(2k)>2(\gamma+1).$
Then Algorithm 3 can be implemented with $\mathcal{O}(k)$ words of memory and $\mathcal{O}(1)$ update time. For any item $x\in\hat{\mathcal{T}}$ with true frequency $f_{x}$, with probability at least $1-\delta$, the released estimate $\hat{C}[x]$ obeys
$f_{x}\;-\;\tfrac{1}{\varepsilon}\ln\tfrac{1}{\delta}\;\leq\;\hat{C}[x]\;\leq\;f_{x}+\frac{|X|}{2k}+\tfrac{1}{\varepsilon}\ln\tfrac{1}{\delta}.$
@@
The proof is relegated to Appendix C.2. Notably, there are inherent accuracy lower bounds for heavy-hitter mechanisms. In the streaming setting, any method that outputs at most $k$ items has a worst-case input on which the error is at least $|X|/(k+1)$ Bose et al. (2003). When differential privacy is imposed, the error must additionally scale as $\mathcal{O}\!\left(\tfrac{1}{\varepsilon}\log\!\left(\tfrac{1}{\delta}\right)\right)$ for some inputs Balcer and Vadhan (2019). Therefore, the error of our mechanism is asymptotically optimal.